# Combinatorial summation of Feynman diagrams: Equation of state of the $2D$ $SU(N)$ Hubbard model

### v3

@@ I Introduction
In the diagrammatic Monte Carlo (DiagMC) approach to correlated systems Prokof’ev and Svistunov (1998, 2007); Van Houcke et al. (2010); Kozik et al. (2010), thermodynamic observables or correlation functions are expressed in terms of a sum of all connected Feynman diagrams of the many-body perturbation theory Abrikosov et al. (2012), which is then sampled stochastically. Each diagram represents a formula for computing one term in this sum, which in the simplest case consists of a product of one-particle non-interacting Green’s functions $G^{0}$ and a number $n$ (the diagram order) of interaction vertices $V$  [^1] For clarity, we confine ourselves to the case of pairwise interactions. (see Fig. 1a,b), integrated over all internal variables. The key advantage of this approach is that the diagrams are defined directly in the thermodynamic limit, circumventing the need to extrapolate the result with the system size, which is typically hard in unbiased quantum Monte Carlo methods due to the negative sign problem Loh et al. (1990); Troyer and Wiese (2005). However, the control of precision in DiagMC relies on its ability to accurately compute the sum of all diagrams to a sufficiently high order $n$, which is inhibited by a factorially increasing with $n$ number of diagrams and hence exploding Monte Carlo variance. Indeed, in correlated regimes all $\sim n!$ order-$n$ diagrams are typically of comparable magnitude Gukelberger et al. (2015), which largely negates the chief advantage of Monte Carlo – that of importance sampling. This suggests that the summation over diagram topologies and indices on which there is only weak dependence could be done deterministically to a similar effect, or more efficiently if such summation could be performed faster than in $\mathcal{O}(n!)$ steps. For fermionic systems, where the diagrams have alternating signs, this also helps lower the Monte Carlo variance Profumo et al. (2015); Rossi (2017); Rossi et al. (2017); Chen and Haule (2019). Crucially, if the computational cost could be reduced to exponential in $n$, it was shown in Ref. Rossi et al. (2017) (with an extension to divergent series  Kim et al. (2021), if necessary) that the computational time would scale only polynomially with the inverse of the desired error bound.
@@
Here we develop a framework for efficient evaluation of Feynman’s diagrammatic series of arbitrary structure by deterministic summation of all diagram integrands. The approach is based on an explicit combinatorial construction of each term in the sum, one Green’s function at a time, whereby at each step the result is maximally factorised into sums of single Green’s functions by dynamic programming. Specifically, the result takes the form of a directed graph (Fig. 1d), with each node being a sum of contributions from its incoming edges, and each edge conveying the value of the previous node multiplied by a Green’s function. In this approach, the $SU(N)$ symmetry is accounted for by merely an additional multiplication of certain edges by a constant factor, while all connected diagrams of order $n$ can be summed in at most $\mathcal{O}(n^{3}4^{n})$ steps independently of $N$. This is reduced for the special case of $N=1$ (spinless fermions) and $SU(2)$ Hubbard model to $\mathcal{O}(n^{3}3^{n})$. The factorisation of the sum, which serves to minimise the number of repeated uses of each Green’s function value, is the essence of the speed-up. As a byproduct, the result is also symmetrised over $n!$ permutations of interaction lines and $2^{n}$ inversions of their end points, helping to further reduce the variance of the DiagMC evaluation of the series. However, the connected topologies with ordered interaction ends can be summed in $\mathcal{O}(n^{3}3^{n})$ steps in the $SU(N)$ case, while those with completely ordered vertices in only $\mathcal{O}(n^{2}2^{n})$ operations. Following Ref. Rossi et al. (2017) (and Kim et al. (2021) in the case of a divergent series), the exponential computational cost of this approach implies polynomial scaling of the calculation time with the required accuracy. The approach admits a vector formulation, which is potentially suitable for a realisation on a quantum computer with a further exponential speed-up.
We apply the combinatorial summation (CoS) technique to a calculation of the equation of state (EoS) of the $2D$ $SU(N)$ Hubbard model in the case of $N=6$, which is relevant for experiments, but hard for numerical methods. We first address the low-temperature regime studied very recently by Pasqualetti et al. Pasqualetti et al. (2024), where the system was realised using the $6$ nuclear spin states of ${}^{173}$b atoms loaded in an optical lattice, and the experimentally obtained EoS was cross-benchmarked against unbiased DQMC calculations. The range of the CoS technique is then explored by extending the calculations to lower temperatures and greater interaction strengths, where the sign problem is known to rapidly intensify Ibarra-García-Padilla et al. (2021) and experimental data for $N=6$ cannot be reliably captured by numerical methods Taie et al. (2022). At the low-temperature/strong-coupling boundary of the studied regime, traits of a developing (pseudo-)gapped state are observed.
II Combinatorial Summation
@@
Here the operators $\hat{c}^{\dagger}_{i\sigma}$ and $\hat{c}_{i\sigma}$ create and annihilate a fermion on site $i$ with the spin $\sigma=1,\ldots,N$, respectively, $\hat{n}_{i\sigma}=\hat{c}^{\dagger}_{i\sigma}\hat{c}_{i\sigma}$ is the number operator, $t$ the hopping amplitude, $U$ the on-site interaction, $\mu$ the chemical potential, and $\langle i,j\rangle$ implies that the summation is over nearest-neighbour lattice sites. A thermodynamic observable, such as, e.g., the average potential energy, is expressed diagrammatically (Fig. 1a,b) as the sum of all connected closed-loop diagrams obtained by linking vertices $\alpha$, representing a point on the lattice $i_{\alpha}$ and in imaginary time $\tau_{\alpha}$, by the interaction lines $V_{\sigma_{\alpha}\sigma^{\prime}_{\alpha}\sigma_{\beta}\sigma^{\prime}_{\beta}}(\alpha,\beta)=\frac{U}{2}\delta_{\sigma_{\alpha},\sigma^{\prime}_{\alpha}}\delta_{\sigma_{\beta},\sigma^{\prime}_{\beta}}(1-\delta_{\sigma_{\alpha},\sigma_{\beta}})\delta_{i_{\alpha},i_{\beta}}\delta(\tau_{\alpha}-\tau_{\beta})$ and non-interacting propagators (Green’s functions) $G^{0}_{\sigma}(\alpha,\beta)=-\langle\mathcal{T}\hat{c}_{i_{\beta}\sigma}(\tau_{\beta})\hat{c}^{\dagger}_{i_{\alpha}\sigma}(\tau_{\alpha})\rangle_{0}$, where $\mathcal{T}$ is the time ordering operator and the statistical average $\langle\ldots\rangle_{0}$ is taken with the Hamiltonian at $U=0$, and summing or integrating the result over all its $\sigma_{\alpha},i_{\alpha},\tau_{\alpha}$ variables. It is well known—and used in finite-size determinant diagrammatic Monte Calro methods Rubtsov et al. (2005); Burovski et al. (2006)—that the sum of all combinations of $n$ interactions with the propagators is generated by the determinant of a $2n\times 2n$ matrix $g_{\alpha\beta}=G^{0}(\alpha,\beta)$, $\alpha,\beta=1,\ldots,2n$ (the spin indices are omitted for clarity), multiplied by the corresponding values of $V(\alpha,\beta)$. This way the $2n!$ terms can be produced extremely efficiently in $\mathcal{O}(n^{3})$ operations, but having to eliminate the unwanted disconnected diagrams from the determinant afterwards requires at least an exponential number of steps Rossi (2017). Our strategy, in contrast, will be to not generate the disconnected diagrams from the start.
II.1 Evaluation of the determinant
@@
A unique element $e$ must appear in $\mathcal{C}$ only once, which in general makes step $l$ dependent on all the steps before it. However, it was demonstrated in Ref. Mahajan and Vinay (1997) that all terms with repeated elements will cancel out due to the sign structure, provided the lowest element in each cycle within $\mathcal{C}$, called the cycle head $h$, is present in $\mathcal{C}$ only once. Then, for each $\mathcal{C}$ with a repeated element, there will be exactly one such $\mathcal{C}^{\prime}$ that $\text{weight}\,\mathcal{C}^{\prime}=\text{weight}\,\mathcal{C}$, but the number of its cycles differs by one, i.e. $\text{sign}\,\mathcal{C}^{\prime}=-\text{sign}\,\mathcal{C}$. This is straightforward to ensure if, at each step $l$, the head of the current cycle $h$ is stored alongside the current element $e$, and the next step is either to any other element $e^{\prime}>h$ within the cycle, or starts a new cycle with $h^{\prime}>h$ and $e^{\prime}=h^{\prime}$. Therefore, each unique node must carry the three numbers $[l,h,e]$, $l=0,\ldots 2n$; $h,e=1,\ldots 2n+1$. The resulting graph, computing the determinant in $\mathcal{O}(n^{4})$ floating-point operations, is illustrated in Fig. 1c.
@@ II.2 Approach 1: modification of the determinant
Since there is a one-to-one correspondence between a particular path in the graph and the diagram it generates, the task of omitting the disconnected diagrams from the determinant can be formulated as that of identifying the corresponding paths and eliminating them selectively. Preserving all other paths is in principle accomplished by duplicating certain nodes along the unwanted paths and re-routing the paths to be kept through the copies, as in the example in Fig. 1d. This suggests that the information $[l,h,e]$, which uniquely identifies the nodes in the determinant, is incomplete for a diagrammatic series obeying more general rules, and the node label must be extended by some additional record $\mathcal{R}$. If what constitutes $\mathcal{R}$ is identified, the right graph can be constructed from the start by the same one-step-at-a-time algorithm, only now the two nodes $[l_{1},h_{1},e_{1}]\otimes\mathcal{R}_{1}$ and $[l_{2},h_{2},e_{2}]\otimes\mathcal{R}_{2}$ are considered identical, and are merged, only if $\mathcal{R}_{1}=\mathcal{R}_{2}$, in addition to $l_{1}=l_{2}$, $h_{1}=h_{2}$, $e_{1}=e_{2}$. In principle, the information in $\mathcal{R}$ should be kept minimal to prevent spawning redundant nodes, but a sub-optimal graph can always be pruned in the end, without changing its value, by merging all nodes with equal $[l,h,e]$ that connect to the same nodes at the next level.
@@
For the $SU(N)$ Hubbard interaction in the form (1), where the same-spin coupling is excluded, the sum over different combinations of spin indices implies that each diagram comes with the spin- and topology-dependent multiplicity factor $M=\sum_{\sigma_{1},\ldots,\sigma_{m}}\prod_{\text{interactions}}(1-\delta_{\sigma_{i},\sigma_{j}})/2$, where $m$ is the number of loops and each interaction in the product connects a loop with spin $\sigma_{i}$ to that with spin $\sigma_{j}$, as in the example in Fig. 1b. A strength of our approach is that an arbitrary factor can be accounted for by merely (i) grouping the diagrams with the same $M$ together and (ii) multiplication of the value of each node at the penultimate level $l=2n-1$ by $M$. To this end, we also store in $\mathcal{R}$ a matrix of connections between the cycles, which does not need to be optimal, and prune the final graph to minimise its size.
Despite the combinatorial structure of $\mathcal{R}$, this algorithm is already efficient at diagram orders typically accessible in calculations. Indeed, since the Monte Carlo variance of integration over the vertex positions and times scales exponentially with diagram order $n$ Rossi et al. (2017), in correlated regimes only contributions from $n\lesssim 10$ can typically be evaluated with $<10\%$ statistical error required for precision reconstruction of observables Šimkovic and Kozik (2019). Fig. 2 shows the actual number of floating point operations required to sum all connected diagrams of order $n$, with this approach labelled there as CoS-1. For the $SU(2)$ Hubbard model, where each diagram has multiplicity $M=1$ and an efficient algorithm (CDet Rossi (2017)) exists, CoS-1 already exhibits competitive performance. In the $SU(N)$ case, the computational cost of CoS-1 is independent of $N$ (for $N>4$) and appears exponential for orders $n\lesssim 6$, although it is expected to eventually rise combinatorially. Nonetheless, at large $n$, CoS-1 is superseded by an approach of exponential complexity described below.
Figure 2: Number of floating-point operations (FLOP) required to evaluate the sum of the integrands of all Feynman diagrams of order $n$: (a) for connected diagrams of the $SU(2)$ Hubbard model summed by the algorithm of Sec. II.2 (CoS-1) and that of Sec. II.3 (CoS-2), for which the curve closely follows $\approx n^{3}3^{n}/8$; the reference dotted line $(2n)^{3}2^{n}+3^{n}$ indicates the theoretical scaling of the CDet algorithm of Ref. Rossi (2017); (b) for connected diagrams in the $SU(N)$ case, with the curve for CoS-2 following $\approx n^{3}4^{n}/7$. Shown as CoS-GW is the computational cost of summing the skeleton (bold-line) series in terms of the renormalised Green’s function $G$ and screened interaction $W$.

### v4

@@ I Introduction
From quantum electro- and chromodynamics to nuclear and condensed matter physics, Feynman’s diagrammatic technique is arguably the most universal approach to describing quantum many-body systems. Scattering amplitudes and decay rates, masses of elementary particles, thermodynamic observables or correlation functions are expressed formally exactly as an infinite sum of all connected Feynman diagrams of the many-body perturbation theory Abrikosov et al. (2012). Each diagram represents a formula for computing one term in this sum, which in the simplest case consists of a product of one-particle non-interacting Green’s functions $G^{0}$ and a number $n$ (the diagram order) of interaction vertices $V$  [^1] For clarity, we confine ourselves to the case of pairwise interactions. (see Fig. 1a,b), integrated over all internal variables. The accuracy of computing a physical observable in a diagrammatic calculation relies on one’s ability to evaluate all terms of the series to a sufficiently high order $n$. This is accomplished by the diagrammatic Monte Carlo (DiagMC) approach Prokof’ev and Svistunov (1998, 2007); Van Houcke et al. (2010); Kozik et al. (2010) through sampling the series stochastically, at the expense of a known and controllable statistical uncertainty. The key advantage of this approach is that the diagrams are defined directly in the thermodynamic limit, circumventing the need to extrapolate the result with the system size, which is typically hard in unbiased quantum Monte Carlo methods due to the negative sign problem Loh et al. (1990); Troyer and Wiese (2005). However, the accuracy of evaluating a high-order contribution is inhibited by a factorially increasing with $n$ number of diagrams and hence exploding Monte Carlo variance. Indeed, in correlated regimes all $\sim n!$ order-$n$ diagrams are typically of comparable magnitude Gukelberger et al. (2015), which largely negates the chief advantage of Monte Carlo – that of importance sampling. This suggests that the summation over diagram topologies and indices on which there is only weak dependence could be done deterministically to a similar effect, or more efficiently if such summation could be performed faster than in $\mathcal{O}(n!)$ steps. For fermionic systems, where the diagrams have alternating signs, this also helps lower the Monte Carlo variance Profumo et al. (2015); Rossi (2017); Rossi et al. (2017); Chen and Haule (2019). Crucially, if the computational cost could be reduced to exponential in $n$, it was shown in Ref. Rossi et al. (2017) (with an extension to divergent series  Kim et al. (2021), if necessary) that the computational time would scale only polynomially with the inverse of the desired error bound.
@@
Here we develop a framework for efficient evaluation of Feynman’s diagrammatic series of arbitrary structure by deterministic summation of all diagram integrands. The approach is based on an explicit combinatorial construction of each term in the sum, one Green’s function at a time, whereby at each step the result is maximally factorised into sums of single Green’s functions by dynamic programming. Specifically, the result takes the form of a directed graph (Fig. 1d), with each node being a sum of contributions from its incoming edges, and each edge conveying the value of the previous node multiplied by a Green’s function. In this approach, the $SU(N)$ symmetry is accounted for by merely an additional multiplication of certain edges by a constant factor, while all connected diagrams of order $n$ can be summed in at most $\mathcal{O}(n^{3}4^{n})$ steps independently of $N$. This is reduced for the special case of $N=1$ (spinless fermions) and $SU(2)$ Hubbard model to $\mathcal{O}(n^{3}3^{n})$. The factorisation of the sum, which serves to minimise the number of repeated uses of each Green’s function value, is the essence of the speed-up. As a byproduct, the result is also symmetrised over $n!$ permutations of the interaction lines and $2^{n}$ inversions of their end points, helping to further reduce the variance of the DiagMC evaluation of the series. However, the connected topologies with ordered interaction ends can be summed in $\mathcal{O}(n^{3}3^{n})$ steps in the $SU(N)$ case, and those with completely ordered vertices in only $\mathcal{O}(n^{2}2^{n})$ operations. Following Ref. Rossi et al. (2017) (and Kim et al. (2021) in the case of a divergent series), the exponential computational cost of this approach implies polynomial scaling of the calculation time with the required accuracy. The approach admits a vector formulation, which is potentially suitable for a realisation on a quantum computer with a further exponential speed-up.
We apply the combinatorial summation (CoS) technique to a calculation of the equation of state (EoS) of the $2D$ $SU(N)$ Hubbard model in the case of $N=6$, which is relevant for experiments, but hard for numerical methods. We first address the low-temperature regime studied very recently by Pasqualetti et al. Pasqualetti et al. (2024), where the system was realised using the $6$ nuclear spin states of 173Yb atoms loaded in an optical lattice, and the experimentally obtained EoS was cross-benchmarked against unbiased DQMC calculations. The range of the CoS technique is then explored by extending the calculations to lower temperatures and greater interaction strengths, where the sign problem is known to rapidly intensify Ibarra-García-Padilla et al. (2021) and experimental data for $N=6$ cannot be reliably captured by numerical methods Taie et al. (2022). At the low-temperature/strong-coupling boundary of the studied regime, traits of a developing (pseudo-)gapped state are observed.
II Results
@@
Here the operators $\hat{c}^{\dagger}_{i\sigma}$ and $\hat{c}_{i\sigma}$ create and annihilate a fermion on site $i$ with the spin $\sigma=1,\ldots,N$, respectively, $\hat{n}_{i\sigma}=\hat{c}^{\dagger}_{i\sigma}\hat{c}_{i\sigma}$ is the number operator, $t$ the hopping amplitude, $U$ the on-site interaction, $\mu$ the chemical potential, and $\langle i,j\rangle$ implies that the summation is over nearest-neighbour lattice sites. A thermodynamic observable, such as, e.g., the average potential energy, is expressed diagrammatically (Fig. 1a,b) as the sum of all connected closed-loop diagrams obtained by linking vertices $\alpha$, representing a point on the lattice $i_{\alpha}$ and in imaginary time $\tau_{\alpha}$, by the interaction lines $V_{\sigma_{\alpha}\sigma^{\prime}_{\alpha}\sigma_{\beta}\sigma^{\prime}_{\beta}}(\alpha,\beta)=\frac{U}{2}\delta_{\sigma_{\alpha},\sigma^{\prime}_{\alpha}}\delta_{\sigma_{\beta},\sigma^{\prime}_{\beta}}(1-\delta_{\sigma_{\alpha},\sigma_{\beta}})\delta_{i_{\alpha},i_{\beta}}\delta(\tau_{\alpha}-\tau_{\beta})$ and non-interacting propagators (Green’s functions) $G^{0}_{\sigma}(\alpha,\beta)=-\langle\mathcal{T}\hat{c}_{i_{\beta}\sigma}(\tau_{\beta})\hat{c}^{\dagger}_{i_{\alpha}\sigma}(\tau_{\alpha})\rangle_{0}$, where $\mathcal{T}$ is the time ordering operator and the statistical average $\langle\ldots\rangle_{0}$ is taken with the Hamiltonian at $U=0$, and summing or integrating the result over all its $\sigma_{\alpha},i_{\alpha},\tau_{\alpha}$ variables. It is well known—and used in finite-size determinant diagrammatic Monte Calro methods Rubtsov et al. (2005); Burovski et al. (2006)—that the sum of all combinations of $n$ interactions with the propagators is generated by the determinant of a $2n\times 2n$ matrix $g_{\alpha\beta}=G^{0}(\alpha,\beta)$, $\alpha,\beta=1,\ldots,2n$ (the spin indices are omitted for clarity), multiplied by the corresponding values of $V(\alpha,\beta)$. This way the $(2n)!$ terms can be produced extremely efficiently in $\mathcal{O}(n^{3})$ operations, but having to eliminate the unwanted disconnected diagrams from the determinant afterwards requires at least an exponential number of steps Rossi (2017). Our strategy, in contrast, will be to not generate the disconnected diagrams from the start.
Figure 2: Number of floating-point operations (FLOP) required to evaluate the sum of the integrands of all Feynman diagrams of order $n$: (a) for connected diagrams of the $SU(2)$ Hubbard model summed by the algorithm of Sec. II.2 (CoS-1) and that of Sec. II.3 (CoS-2), for which the curve closely follows $\approx n^{3}3^{n}/8$; the reference dotted line corresponds to the state-of-the-art implementation of the CDet algorithm Rossi (2017) with fast principal minor calculation IV and Ferrero (2022) ($\sum_{l=0}^{n-2}2^{l+1}(n-l-1)^{2}+3^{n}$); (b) for connected diagrams in the $SU(N)$ case, with the curve for CoS-2 following $\approx n^{3}4^{n}/7$. Shown as CoS-GW is the computational cost of summing the skeleton (bold-line) series in terms of the renormalised Green’s function $G$ and screened interaction $W$. Source data are provided as a Source Data file.
II.1 Combinatorial Summation for the determinant
@@
A unique element $e$ must appear in $\mathcal{C}$ only once, which in general makes step $l$ dependent on all the steps before it. However, it was demonstrated in Ref. Mahajan and Vinay (1997) that all terms with repeated elements will cancel out due to the sign structure, provided the lowest element in each cycle within $\mathcal{C}$, called the cycle head $h$, is placed at the beginning of the cycle and is present in $\mathcal{C}$ only once. Then, for each $\mathcal{C}$ with a repeated element, there will be exactly one $\mathcal{C}^{\prime}$ such that $\text{weight}\,\mathcal{C}^{\prime}=\text{weight}\,\mathcal{C}$, but the number of its cycles differs by one, i.e. $\text{sign}\,\mathcal{C}^{\prime}=-\text{sign}\,\mathcal{C}$. This is straightforward to ensure if, at each step $l$, the head of the current cycle $h$ is stored alongside the current element $e$, and the next step is either to any other element $e^{\prime}>h$ within the cycle, or starts a new cycle with $h^{\prime}>h$ and $e^{\prime}=h^{\prime}$, so that the cycle heads are picked in ascending order. Therefore, each unique node must carry the three numbers $[l,h,e]$, $l=0,\ldots 2n$; $h,e=1,\ldots 2n+1$. The resulting graph, computing the determinant in $\mathcal{O}(n^{4})$ floating-point operations Mahajan and Vinay (1997), is illustrated in Fig. 1c.
@@ II.2 Approach 1: modification of the determinant
Since there is a one-to-one correspondence between a particular path in the graph and the diagram it generates, the task of omitting the disconnected diagrams from the determinant can be formulated as that of identifying the corresponding paths and eliminating them selectively. Preserving all other paths is in principle accomplished by duplicating certain nodes along the unwanted paths and re-routing the paths to be kept through the copies, as in the example in Fig. 1d. This suggests that the information $[l,h,e]$, which uniquely identifies the nodes in the determinant, is incomplete for a diagrammatic series obeying more general rules, and the node label must be extended by some additional record $\mathcal{R}$. If what constitutes $\mathcal{R}$ is identified, the right graph can be constructed from the start by the algorithm of Sec. II.1 with the modification that the two nodes $[l_{1},h_{1},e_{1}]\otimes\mathcal{R}_{1}$ and $[l_{2},h_{2},e_{2}]\otimes\mathcal{R}_{2}$ are considered one and the same if $\mathcal{R}_{1}=\mathcal{R}_{2}$ in addition to $l_{1}=l_{2}$, $h_{1}=h_{2}$, $e_{1}=e_{2}$. It is desirable that the information in $\mathcal{R}$ is kept minimal to prevent spawning redundant nodes, but a sub-optimal graph can always be pruned in the end, without changing its value, by merging all nodes with equal $[l,h,e]$ that connect to the same nodes at the next level.
@@
For the $SU(N)$ Hubbard interaction in the form (1), where the same-spin coupling is excluded, the sum over different combinations of spin indices implies that each diagram comes with the spin- and topology-dependent multiplicity factor $M=\sum_{\sigma_{1},\ldots,\sigma_{m}}\prod_{\text{interactions}}(1-\delta_{\sigma_{i},\sigma_{j}})/2$, where $m$ is the number of loops and each interaction in the product connects a loop with spin $\sigma_{i}$ to that with spin $\sigma_{j}$, as in the example in Fig. 1b. A strength of our approach is that an arbitrary factor can be accounted for by merely (i) grouping the diagrams with the same $M$ together and (ii) multiplication of the value of each node at the penultimate level $l=2n-1$ by $M$. To this end, we also store in $\mathcal{R}$ a matrix of connections between the cycles and prune the final graph to minimise its size.
Despite the combinatorial structure of $\mathcal{R}$, this algorithm is already efficient at diagram orders typically accessible in calculations. Indeed, since the Monte Carlo variance of integration over the vertex positions and times scales exponentially with diagram order $n$ Rossi et al. (2017), in correlated regimes only contributions from $n\lesssim 10$ can typically be evaluated with $<10\%$ statistical error required for precision reconstruction of observables Šimkovic and Kozik (2019). Fig. 2 shows the actual number of floating point operations required to sum all connected diagrams of order $n$, with this approach labelled there as CoS-1. For the $SU(2)$ Hubbard model, where each diagram has multiplicity $M=1$ and an efficient algorithm exists—connected determinant diagrammatic Monte Carlo (CDet) Rossi (2017) with the fast principal minor algorithm IV and Ferrero (2022)—CoS-1 already exhibits competitive performance. In the $SU(N)$ case, the computational cost of CoS-1 is independent of $N$ (for $N>4$) and appears exponential for orders $n\lesssim 6$, although it is expected to eventually rise combinatorially. Nonetheless, at large $n$, CoS-1 is superseded by an approach of exponential complexity described below.